OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable. y2/y - 3 . y2 - y - 6/ y2 + 1y

OpenStudy (jack1):

\[\large \frac {y^2}{y - 3} \times \frac {y^2 - y - 6}{ y^2 + y}\] is this your equation @MsJasmine1 ?

OpenStudy (jack1):

\[\large \frac {y^2}{y - 3} \times \frac {(y - 3)(y + 2) }{ y^2 + y}\] \[\large =\frac {(y^2)(y - 3)(y + 2) }{ (y - 3)(y^2 + y)}\]

OpenStudy (jack1):

can you simplify from here?